OpenStudy (anonymous):

What conclusion can be based on the given statement? y is a number such that 6y = 12 y=1/2 y=2 y=72

OpenStudy (texaschic101):

6y = 12 (divide both sides by 6) 6y/6 = 12/6 y = 2